OpenStudy (johnnydicamillo):

find the limit when x approaches 0

OpenStudy (johnnydicamillo):

\[\lim_{x \rightarrow 0}\frac{ \sqrt{9 + h} - 3 }{ h }\]

OpenStudy (johnnydicamillo):

I know I need to rationalize the numerator but I think I am doing some mathematical error. Can someone show me how to rationalize the numerator?

geerky42 (geerky42):

You mean \[\large \lim_{h\to0}\dfrac{\sqrt{9+h}-3}{h}\] right? You should rationalize numerator, then simplify it. Works for me. What went wrong?

geerky42 (geerky42):

To rationalize numeration, you just multiply numerator and denominator by conjugate of numerator: \(\sqrt{9+h}+3\)

geerky42 (geerky42):

numerator *

OpenStudy (johnnydicamillo):

and I get \[\frac{ 9 + h - 9 }{ h \sqrt{9 + h} + 3 }\]

OpenStudy (johnnydicamillo):

but it the numerator would still end up being 0

geerky42 (geerky42):

Try to simplify.

OpenStudy (freckles):

and the sqrt(9+h)+3 should have ( ) around it (on bottom)

OpenStudy (johnnydicamillo):

okay, um... give me a sec

OpenStudy (johnnydicamillo):

i am stuck

geerky42 (geerky42):

You have \(\dfrac{9+h-9}{h(\sqrt{9+h}+3)}\), right? \[\dfrac{9+h-9}{h(\sqrt{9+h}+3)}\\~\\\dfrac{h}{h(\sqrt{9+h}+3)}\\~\\\dfrac{\cancel h}{\cancel h(\sqrt{9+h}+3)}\\~\\\dfrac{1}{\sqrt{9+h}+3}\]

OpenStudy (johnnydicamillo):

got, it.. For some reason the h just disappeared in my head. Thanks

geerky42 (geerky42):

No problem